OpenStudy (anonymous):

Find the derivative of f(x) = -12x2 + 9x at x = 6.

OpenStudy (anonymous):

\[\frac{ d }{ dx }(-12x ^{2}+9x)=(2)(-12)x ^{2-1}+(1)9x ^{1-1}\] \[=-24x+9|_{6}=(-24)(6)+9=-144+9=-135\]

OpenStudy (anonymous):

derivatives follow the form \[nx ^{n-1}=\frac{ dy }{ dx }\]